a plate of mass m is held at rest by firing bullets from below. each bullet has a mass m, velocity u(up) just before hitting and stops after hitting the platefor a moment and falls. determine the number of bullets striking the plate per unit time?

Grade:11

1 Answers

Batman
7 years ago
Force Down = Force Up
m * g = Rate Of Change of Momentum up
m * g = Rate of bullets * m * u
Rate of bullets = g/u (g=gravity)